a car accelerate from rest to reach a certain speed in 10 min. it then continues at this speed for another 10 minutes and decelerate to rest in another 5 min, the total distance covered is 17.5km. find the steady speed reached


1
Expert's answer
2021-04-18T19:32:43-0400

Let the times of motion during three different parts of the distance be "t_1, t_2, t_3", the constant speed after the acceleration "v_1", and accelerations during first and third intervals be "a_1, a_3". Then the corresponding displacements and speeds during three parts of the way are:

1) "s_1 = \frac{a_1 t_1^2}{2}", "v_1 = a_1 t_1"

2) "s_2 = v_1 t_2"

3) "s_3 = v_1 t_3 - \frac{a_3 t_3^2}{2}, 0 = v_1 - a_3 t_3".

Total distance is "s = s_1 + s_2 + s_3 = \frac{a_1 t_1^2}{2} + v_1 (t_2 + t_3) - \frac{a_3 t_3^2}{2}".

From 1) and 3), "v_1 = a_1 t_1 = a_3 t_3", from where "a_3 = \frac{t_1}{t_3} a_1". Substituting the last expression for "a_3" and "v_1 = a_1 t_1" from 1) into "s", obtain:

"s = \frac{a_1 t_1^2}{2} + a_1 t_1(t_2 + t_3) - \frac{a_1 t_1 t_3}{2}", from where "a_1 = \frac{s}{\frac{t_1^2}{2} + t_1 t_2 + \frac{t_1 t_3}{2}} = 0.1 \frac{km}{min^2}".

The steady speed is hence "v_1 = a_1 t_1 = 0.1 \frac{km}{min^2} \cdot 10 min = 1\frac{km}{min}".